# Quantum control of continuous systems via nonharmonic potential modulation

### v3

@@ 1 Introduction
The preparation of a continuous-variable system in a non-Gaussian quantum state is of paramount importance in various aspects of quantum science. This ranges from fundamental tests of quantum mechanics [1, 2, 3, 4, 5], through the design of quantum sensors [6, 7, 8, 9], to quantum information processing [10, 11, 12, 13, 14, 15, 16, 17, 18]. The generation of non-Gaussian states requires a nonlinear resource, often introduced through coupling to an auxiliary degree of freedom, e.g., a two-level system [19, 20, 12, 21, 15, 22, 5]. On the other hand, some continuous-variable systems already possess intrinsic nonharmonicity in the potential of a canonical variable. Notable examples include the motion of a par-
Figure 1: (a) Examples of continuous-variable nonlinear systems that can be optimally controlled without the need of auxiliary systems—single atoms in optical tweezers and flux-tunable transmons. (b) Time-evolved probability density $P(x,\tau)=|\psi(x,\tau)|^{2}$, during the state preparation protocol with the snapshots of Wigner functions $W(x,p)$ at the beginning and the end of the protocol. The potential is Gaussian and its optimally controlled position $u(\tau)$ is depicted via a solid black line. The top panel shows a comparison between a weak and long sinusoidal drive resonant with the ground-first excited state transition and an optimized much faster control. The bottom panel presents an optimal control leading to the GKP state. (c) Excited states occupation numbers $|c_{n}(\tau)|^{2}=|\innerproduct{\psi_{n}(x)}{\psi(x,\tau)}|^{2}$ for $\psi_{0}(x)\rightarrow\psi_{1}(x)$ Fock excitation protocols of (b)top. The left panel corresponds to the slow Rabi flop, while the right one shows the fast optimal control. Utilizing more states within the nonharmonic potential accelerates and makes the control more versatile.
@@
ticle in a trap of finite depth [23, 24] and the current in an electric circuit with a Josephson junction [25]. These nonharmonicites in the potential are typically used to define a qubit within continuous-variable systems [26, 27, 28, 29, 30]. In contrast, we explore methods for utilizing this intrinsic nonlinearity to generate and control states beyond the two-dimensional subspace.
@@
More specifically, we develop protocols to generate a plethora of states, including Fock, Gottesman-Kitaev-Preskill (GKP) [10], multi-legged-cat [1, 31], and cubic-phase [10, 32], using optimal control [33, 34, 35] of the potential’s position and depth. Furthermore, we employ this control mechanism to design protocols that implement arbitrary unitaries in selected subspaces, enable single-shot orthogonal state discrimination [36, 37], spatial state transfer, correct for nonlinear evolution, and perform algorithmic cooling [38, 39]. These protocols could be implemented with single atoms in optical tweezers [40, 41, 42, 43, 44] and lattices [23, 45] to extend the motional states available in protocols with itinerant particles, e.g., fermionic quantum processors [18]; or with flux-tunable transmons [28, 46, 25] for minimally invasive state manipulation [cf. Fig 1(a)]. Previous works, including control of Bose-Einstein condensates [47, 48, 49, 50, 51, 52, 53, 54, 55, 56], fast atom transport [57, 58, 45], ion shuttling [59], cat state creation through two-photon driving [60], and lattice interferometry [61] have demonstrated similar concepts. However, the impact of nonharmonic potentials in quantum information processing has yet to be fully realized, despite being a natural area of exploration—especially given the potential to utilize, for example, the motion of single atoms in scalable optical tweezer arrays [39].
@@
The paper is structured as follows. Section 2 introduces details about the physical model and optimization techniques, while Section 3 presents several protocols that can be achieved through potential modulation, including state preparation, implementation of unitaries, single-shot measurements, cooling, correcting for nonlinear evolution, and spatial state transfer. Further, Section 4 analyzes the feasibility of our proposal, addressing the speed limits of the protocols, their robustness against decoherence, and how couplings to other modes can be incorporated into optimization. Section 5 concludes the paper with a finishing discussion and an outlook.
@@ 2.1 Hamiltonians for different systems
where $m$ is the mass of an atom, $\hat{X}$ is the position operator, $\hat{P}$ is the momentum operator, $t$ is time, and functions $a(t)$ and $U(t)$ control the depth and position of the potential. Here, the trapping potential $V(\hat{X})$ is Gaussian for the tweezer,
@@
Here, $\hat{\varphi}=(\hat{\varphi}_{1}+\hat{\varphi}_{2})/2$ is the phase difference operator, $\hat{n}=\hat{Q}/2e$ is charge number operator, $\hat{\varphi}_{i}=2\pi\hat{\Phi}_{i}/\Phi_{0}$ is phase difference across $i^{\text{th}}$ junction, $\Phi_{0}=h/2e$ is the flux quantum, $E_{C}/h={e^{2}}/{2hC}$ is the charge energy, $C$ is the total capacitance, $E_{\text{J}}=E_{\text{J}1}+E_{\text{J}2}$ is the total Josephson energy, $E_{\text{J}i}$ is the Josephson energy of $i^{\text{th}}$ junction, $d=(E_{\text{J}2}-E_{\text{J}1})/(E_{\text{J}2}+E_{\text{J}1})$ is the junction asymmetry, and $\phi_{\text{x}}(t)$ is the external flux through SQUID.
@@
$\hat{x}$ $=\frac{\hat{X}}{X_{0}}=\frac{\hat{n}}{n_{0}}=\frac{1}{\sqrt{2}}(\hat{a}^{\dagger}+\hat{a}),$
$\hat{p}$ $=\frac{\hat{P}}{P_{0}}=\frac{\hat{\varphi}}{\varphi_{0}}=\frac{1}{\sqrt{2}}i(\hat{a}^{\dagger}-\hat{a}),$ (5)
where $\hat{a}^{\dagger}$ creates a single excitation. For an atom trapped in a tweezer, we have
@@
where $v=V/\hbar\omega$ is the dimensionless potential. The control of the system is performed through optimal modulation of position, $u(\tau)$, and depth, $a(\tau)$, of the potential. While for an atom in an optical tweezer, they are independent controls realized through, e.g., acousto-optic modulator [77], for flux-tunable transmons they are constrained through a single control function, the intensity of the external flux [79],
@@
Such modulations have been realized in various other experimental platforms [81, 61, 45, 54, 55], where the choice of either position or depth control depends on feasibility within a specific setup. For example, Paul traps allow for both position and depth control via varying current through electrodes generating the electric field [54], while in optical lattices phase control allows very precise position control [45]. As for optical tweezers, the use of multiple radiofrequency tones makes a powerful knob for the time-dependent control of both position and depth [44].
@@
when expanded up to the second order in the small parameter $\eta$. For the examples considered in this work, the tweezer potential then reads as
@@
Note we have defined the units so that the functional forms of these potentials match up to the second leading order, allowing for a unified discussion of nonharmonicity. Here, $\eta$ is the measure of the nonharmonicity of the potential, comparing the canonical length to the characteristic length scale of the potential, and translates directly into an energy level splitting,
@@
where $\omega_{ij}$ is a transition frequency between $i^{\text{th}}$ and $j^{\text{th}}$ levels. For instance, for an atom in a tweezer it is
@@
for the flux-tunable transmon it is a monotonic function of transmon anharmonicity $\alpha$,
@@
and for Kerr oscillators, it can be associated with Kerr nonlinearity $\kappa\sim\omega\eta^{2}/24$ [82]. The typical values of physical parameters for the examples introduced in this section lead to nonharmoncities no larger than $\eta\sim 0.5$ (see Tab. 1) and can be tuned down at least an order of magnitude depending on the specific setup. Hence, here we will analyze such a parameter regime. Nevertheless, many other experimental platforms are characterized by much lower values, such as trapped ions [83, 24] or levitated mechanical oscillators [84, 85]. There, lower nonharmonicity is associated with either larger potential length scales or larger masses of trapped objects. In such cases, the additional enhancement of nonharmonicity is needed, e.g., through state excitation [86, 87, 85], to generate non-Gaussian states.
@@ 2.2 Control optimization
Before presenting specific protocols, let us discuss possible methods of designing controls $u(\tau)$ and $a(\tau)$. Within each of the subsequent protocols, we aim to achieve some specific goal—be it state preparation, unitary implementation, or others—that can be quantified through a reward function that ought to be maximized. A set of methods for designing time-varying controls for the maximization of such a reward function is collectively called quantum optimal control (QOC) [88, 33, 34]. Introduced in the eighties and since then broadly developed, QOC has become one of the main quantum control tools, among other optimization approaches, such as adiabatic passages [89], shortcuts to adiabaticity [90], and composite pulses [91]. Various techniques have been established to design the control pulses, relying on both gradient-free and gradient-based approaches. The latter assume the ability to differentiate the reward function and include, among the others, GRAPE (gradient ascent pulse engineering)- [92] and Krotov-like [93, 94] techniques for optimizing controls that are piecewise constant in time. On the other hand, the former rely only on the evaluation of the reward function itself. The examples are numerous, including Nelder-Mead [95], evolution strategies [96], simulated annealing [97], and many others. Notably, recent years have brought a rapid surge in using machine learning approaches for quantum control problems [98]. The choice of approach relies on constraints given by a specific experimental setup, including how accurately and fast we can solve the dynamics and what the limitations of control pulses are, such as maximal Fourier bandwidth and intensity. We choose to utilize the dressed chopped random-basis technique (dCRAB) [99] that is based on the Nelder-Mead gradient-free method. Here, the control function is expanded into a Fourier basis with randomized frequencies and a high-frequency cutoff corresponding to approximations of experimentally accessible bandwidths, i.e.,
@@
where $a_{k}$ and $b_{k}$ are parameters to be optimized, $\nu_{k}$ are frequencies that are probabilistically drawn from the uniform distribution in the range $[0,\nu_{\text{max}}]$, $\nu_{\text{max}}$ is the frequency cutoff, and $N_{\text{p}}$ is the number of frequency components. We have chosen this particular gradient-free method as we consider several different reward functions, it naturally implements a frequency cutoff for the control function, and there are available well-developed open-source packages, including the Quantum Optimal Control Suite [35] which we utilize for the optimization. However, depending on the specifics of the experimental implementation, our results can be achieved with an alternative approach, such as, e.g., GRAPE.
For all the simulations, the split-step method [100] was used to simulate the dynamics on a spatial grid. In the case of one-dimensional calculations, throughout optimization runs, the numerical parameters read: number of spatial grid points $N_{x}=2^{8}$, where the grid spanned $x\in[-13,13]$, and number of time grid points $N_{\tau}=500$. After optimizing the pulses, the dynamics was run and further optimized with increased accuracy to make sure the results are converged. The optimization has been performed with QuoCS library version 0.0.44 [35]. The number of Fourier components used for dCRAB varied from 20 to 50 depending on a particular protocol and were randomized from a uniform distribution on a bandwidth that also varied between different protocols. The choice of these optimization parameters was fine-tuned for each of the protocols, however, it was not exclusive—different choices usually yielded similar values of specific reward functions, but with different total optimization iterations. Note that further optimization of all the protocols towards lower infidelities is possible and would involve further fine-tuning of optimization parameters.
@@ 3 Protocols
In the following section, we discuss several protocols that can be implemented via a nonharmonic potential modulation. It includes a versatile non-Gaussian state preparation, both in a single- (sec. 3.1) and double-well (sec. 3.3) potential landscape, as well as the implementation of unitaries (sec. 3.2), single-shot measurements (sec. 3.4), cooling (sec. 3.5), and correcting state evolution for nonlinear evolution due to nonharmonicities (sec. 3.6).
@@ 3.1 State preparation
$\mathcal{F}=\left|\innerproduct{\psi(x,\tau=\tau_{\text{max}})}{\psi_{\text{T}}(x)}\right|^{2}$ (21)
with the target state $\psi_{\text{T}}(x)$. We consider the preparation of several states, including Fock, finite GKP, cubic-phase, and cat states with high fidelity in a one-dimensional geometry. As these states are non-Gaussian, the role of nonharmonicity of the potential is evident—if it was quadratic, and dynamics linear, the Gaussianity of the state would be preserved during the evolution. In Fig. 1(b) we present an example of the excitation protocol, where we specialize to a Gaussian potential with $\eta=0.25$. The target state is a finite GKP state,
@@
Figure 2: Examples of a state preparation protocol involving Gaussian [(a), (c), (e), (g)] and cosine [(b), (d), (f), (h)] single-well potentials, with $\eta=0.25$ and $\tau_{\text{max}}/2\pi=6$, using only optimized position of the potential $u(\tau)$ in the case of the Gaussian potential. The protocols include second excited Fock [(a-b)], cat [(c-d), $s=9$], cubic-phase [(e-f), $\kappa=2$, $r=0.7$], and GKP [(g-h), $r=0.6$, $(d_{1},d_{2},d_{3})=(-\sqrt{4\pi},0,\sqrt{4\pi})$, $(h_{1},h_{2},h_{3})=(1,2,1)$] state preparation. Protocol fidelities $\mathcal{F}$ are: (a) $99.8\%$ (b) $99.8\%$ (c) $92.7\%$ (d) $93.4\%$ (e) $97.2\%$ (f) $94.9\%$ (g) $99.8\%$ (h) $99.8\%$.

### v4

@@ 1 Introduction
The preparation of a continuous-variable system in a non-Gaussian quantum state is of paramount importance in various aspects of quantum science. This ranges from fundamental tests of quantum mechanics [1, 2, 3, 4, 5], through the design of quantum sensors [6, 7, 8, 9], to quantum information processing [10, 11, 12, 13, 14, 15, 16, 17, 18]. The generation of non-Gaussian states requires a nonlinear resource, often introduced through coupling to an auxiliary degree of freedom, e.g., a two-level system [19, 20, 12, 21, 15, 22, 5]. On the other hand, some continuous-variable systems already possess intrinsic nonharmonicity in the potential of a canonical variable (see Fig. 1). Notable examples include the mo-
[…]
@@
tion of a particle in a trap of finite depth [23, 24] and the current in an electric circuit with a Josephson junction [25]. These nonharmonicities in the potential are typically used to define a qubit within continuous-variable systems [26, 27, 28, 29, 30]. In contrast, we investigate whether this intrinsic nonlinearity suffices to implement quantum protocols beyond the two-dimensional subspace. Since it is weaker than the nonlinearity provided by auxiliary two-level systems, we aim to assess how far optimal control can compensate for this limitation.
@@
More specifically, we develop protocols to generate a plethora of states, including Fock, Gottesman-Kitaev-Preskill (GKP) [10], multi-legged-cat [1, 31], and cubic-phase [10, 32], using optimal control [33, 34, 35] of the potential’s position and depth. Furthermore, we employ this control mechanism to design protocols that implement arbitrary unitaries in selected subspaces, enable single-shot orthogonal state discrimination [36, 37], spatial state transfer, correct for non-Gaussian evolution, and perform algorithmic cooling [38, 39]. These protocols could be implemented with single atoms in optical tweezers [40, 41, 42, 43, 44] and lattices [23, 45] to extend the motional states available in protocols with itinerant particles, e.g., fermionic quantum processors [18]; or with flux-tunable transmons [28, 46, 25] for minimally invasive state manipulation [cf. Fig. 1(a)]. Previous works, including control of Bose-Einstein condensates [47, 48, 49, 50, 51, 52, 53, 54, 55, 56], fast atom transport [57, 58, 45], ion shuttling [59], cat state creation through two-photon driving [60], and lattice interferometry [61] have demonstrated similar concepts. However, the role of nonharmonic potentials in quantum information processing remains underexplored, despite their natural relevance—particularly for exploiting, for example, the motion of single atoms in scalable optical tweezer arrays [39].
@@
The paper is structured as follows. Section 2 introduces details about the physical model and optimization techniques, while Section 3 presents several protocols that can be achieved through potential modulation, including state preparation, correcting for non-Gaussian evolution, implementation of unitaries, single-shot measurements, cooling, and spatial state transfer. Further, Section 4 analyzes the feasibility of our proposal, addressing the speed limits of the protocols, their robustness against decoherence, and how couplings to other modes can be incorporated into optimization. Section 5 concludes the paper with a finishing discussion and an outlook.
@@ 2.1 Hamiltonians for different systems
where $m$ is the mass of an atom, $\hat{X}$ is the position operator, $\hat{P}$ is the momentum operator, $t$ is time, and functions $a(t)$ and $U(t)$ control the depth and position of the potential. Here, the trapping potential V(X^)V($\hat${X}) is Gaussian for the tweezer,
@@
Here, the phase difference operator is
$\hat{\varphi}=\frac{\hat{\varphi}_{1}+\hat{\varphi}_{2}}{2},$ (5)
where $\hat{\varphi}_{i}=\frac{2\pi\hat{\Phi}_{i}}{\Phi_{0}}$ denotes the phase difference across the $i^{\text{th}}$ Josephson junction, and $\Phi_{0}=h/2e$ is the magnetic flux quantum. The charge number operator is defined as
$\hat{n}=\frac{\hat{Q}}{2e},$ (6)
and the charge energy is
$\frac{E_{C}}{h}=\frac{e^{2}}{2hC},$ (7)
where $C$ is the total capacitance. The total Josephson energy is
$E_{\text{J}}=E_{\text{J}1}+E_{\text{J}2},$ (8)
with $E_{\text{J}i}$ representing the Josephson energy of the $i^{\text{th}}$ junction. The asymmetry between the two junctions is characterized by
$d=\frac{E_{\text{J}2}-E_{\text{J}1}}{E_{\text{J}2}+E_{\text{J}1}}.$ (9)
Finally, $\phi_{\text{x}}(t)$ denotes the time-dependent external magnetic flux threading the SQUID loop.
@@
[…]
where $\hat{a}^{\dagger}$ creates a single excitation and $\hat{x},\hat{p}]=i$. For an atom trapped in a tweezer, we have
@@
where $v=V/\hbar\omega$ is the dimensionless potential. The control of the system is performed through optimal modulation of position, $u(\tau)$, and depth, $a(\tau)$, of the potential. While for an atom in an optical tweezer, they are independent controls realized through, e.g., acousto-optic modulator [77], for flux-tunable transmons, they are constrained through a single control function, the intensity of the external flux [79],
@@
Such modulations have been realized in various other experimental platforms [81, 61, 45, 54, 55], where the choice of either position or depth control depends on feasibility within a specific setup. For example, Paul traps allow for both position and depth control via varying current through electrodes generating the electric field [54], while in optical lattices, phase control allows very precise position control [45]. As for optical tweezers, the use of multiple radiofrequency tones makes a powerful knob for the time-dependent control of both position and depth [44].
@@
when expanded up to the second order in the small parameter $\eta$. Note that, as the potential $v(x)$ is beyond quadratic in $x$, the ensuing dynamics is nonlinear, i.e., it does not preserve the Gaussianity of the states. For the examples considered in this work, the tweezer potential then reads as
@@
Note that we have defined the units so that the functional forms of these potentials match up to the second leading order, allowing for a unified discussion of nonharmonicity. Here, $\eta$ is the measure of the nonharmonicity of the potential, comparing the canonical length to the characteristic length scale of the potential, and translates directly into an energy level splitting,
@@
where $\omega_{ij}$ is a transition frequency between $i^{\text{th}}$ and $j^{\text{th}}$ levels. For instance, for an atom in a tweezer, it is
@@
for the flux-tunable transmon, it is a monotonic function of transmon anharmonicity $\alpha$,
@@
and for Kerr oscillators, it can be associated with Kerr nonlinearity $\kappa\sim\omega\eta^{2}/24$ [82]. The typical values of physical parameters for the examples introduced in this section lead to nonharmonicities no larger than $\eta\sim 0.5$ (see Tab. 1) and can be tuned down at least an order of magnitude depending on the specific setup. Hence, here we will analyze such a parameter regime. Nevertheless, many other experimental platforms are characterized by much lower values, such as trapped ions [83, 24] or levitated mechanical oscillators [84, 85]. There, lower nonharmonicity is associated with either larger potential length scales or larger masses of trapped objects. In such cases, the additional enhancement of nonharmonicity is needed, e.g., through state excitation [86, 87, 85], to generate non-Gaussian states.
@@ 2.2 Control optimization
Before presenting specific protocols, let us discuss possible methods of designing controls $u(\tau)$ and $a(\tau)$. Within each of the subsequent protocols, we aim to achieve some specific goal—be it state preparation, unitary implementation, or others—that can be quantified through a reward function that ought to be maximized. A set of methods for designing time-varying controls for the maximization of such a reward function is collectively called quantum optimal control (QOC) [88, 33, 34]. Introduced in the eighties and since then broadly developed, QOC has become one of the main quantum control tools, among other optimization approaches, such as adiabatic passages [89], shortcuts to adiabaticity [90], and composite pulses [91]. Various techniques have been established to design the control pulses, relying on both gradient-free and gradient-based approaches. The latter assume the ability to differentiate the reward function and include, among others, GRAPE (gradient ascent pulse engineering)- [92] and Krotov-like [93, 94] techniques for optimizing controls that are piecewise constant in time. On the other hand, the former rely only on the evaluation of the reward function itself. The examples are numerous, including Nelder-Mead [95], evolution strategies [96], simulated annealing [97], and many others. Notably, recent years have brought a rapid surge in the use of machine learning approaches for quantum control problems [98]. The choice of approach relies on constraints given by a specific experimental setup, including how accurately and fast we can solve the dynamics and what the limitations of control pulses are, such as maximal Fourier bandwidth and intensity. We choose to utilize the dressed chopped random-basis technique (dCRAB) [99] that is based on the Nelder-Mead gradient-free method. Here, the control function is expanded into a Fourier basis with randomized frequencies and a high-frequency cutoff corresponding to approximations of experimentally accessible bandwidths, i.e.,
@@
where $a_{k}$ and $b_{k}$ are parameters to be optimized, $\nu_{k}$ are frequencies that are probabilistically drawn from the uniform distribution in the range $0,\nu_{\text{max}}] [$100], $\nu_{\text{max}}$ is the frequency cutoff, and $N_{\text{p}}$ is the number of frequency components. We have chosen this particular gradient-free method as we consider several different reward functions, it naturally implements a frequency cutoff for the control function, and there are available well-developed open-source packages, including the Quantum Optimal Control Suite [35], which we utilize for the optimization. However, depending on the specifics of the experimental implementation, our results can be achieved with an alternative approach, such as, e.g., GRAPE.
Generally speaking, controllability of the system described by Eq. (1) depends sensitively on the shape and depth of the potential. For shallow potentials, the finite number of bound states can limit the set of achievable state transformations. However, in the case of sufficiently deep or unbounded potentials—such as the quartic potential—full controllability can be established under certain control schemes [101]. The chosen potential landscape is then crucial in determining the accessible Hilbert space and, consequently, the effectiveness of optimal control protocols.
For all the simulations, the split-step method [102] was used to simulate the dynamics on a spatial grid. In the case of one-dimensional calculations, throughout optimization runs, the numerical parameters read: number of spatial grid points $N_{x}=2^{8}$, where the grid spanned $x\in[-13,13]$, and number of time grid points $N_{\tau}=500$. After optimizing the pulses, the dynamics was run and further optimized with increased accuracy to make sure the results are converged. The optimization has been performed with QuoCS library version 0.0.44 [35]. The number of Fourier components used for dCRAB varied from 20 to 50, depending on a particular protocol, and was randomized from a uniform distribution on a bandwidth that also varied between different protocols. The choice of these optimization parameters was fine-tuned for each of the protocols, however, it was not exclusive—different choices usually yielded similar values of specific reward functions, but with different total optimization iterations. Note that further optimization of all the protocols towards lower infidelities is possible and would involve further fine-tuning of optimization parameters.
@@ 3 Protocols
In the following section, we discuss several protocols that can be implemented via a nonharmonic potential modulation. It includes a versatile Gaussian (Sec. 3.2) and non-Gaussian state preparation, both in a single- (Sec. 3.1) and double-well (Sec. 3.5) potential landscape, as well as the implementation of unitaries (Sec. 3.4), single-shot measurements (Sec. 3.6), cooling (Sec. 3.7), and correcting state evolution for nonlinear evolution due to nonharmonicities (Sec. 3.3).
Figure 2: Examples of a state preparation protocol involving Gaussian [(a), (c), (e), (g)] and cosine [(b), (d), (f), (h)] single-well potentials, with $\eta=0.25$ and $\tau_{\text{max}}/2\pi=6$, using only optimized position of the potential $u(\tau)$ in the case of the Gaussian potential. The protocols include second excited Fock [(a-b)], cat [(c-d), $s=9$], cubic-phase [(e-f), $\kappa=2$, $r=0.7$], and GKP [(g-h), $r=0.6$, $(d_{1},d_{2},d_{3})=(-\sqrt{4\pi},0,\sqrt{4\pi})$, $(h_{1},h_{2},h_{3})=(1,2,1)$] state preparation. Protocol fidelities $\mathcal{F}$ are: (a) $99.8\%$ (b) $99.8\%$ (c) $92.7\%$ (d) $93.4\%$ (e) $97.2\%$ (f) $94.9\%$ (g) $99.8\%$ (h) $99.8\%$. See the caption of Fig. 1 for subplot and curve legend details.
@@ 3.1 State preparation
[…]
with the target state $\psi_{\text{T}}(x)$. We consider the preparation of several states, including Fock, finite GKP, cubic-phase, and cat states with high fidelity in a one-dimensional geometry. As these states are non-Gaussian, the role of nonharmonicity of the potential is evident—if it was quadratic, and dynamics linear, the Gaussianity of the state would be preserved during the evolution. In Fig. 1(b), we present an example of the excitation protocol, where we specialize to a Gaussian potential with $\eta=0.25$. The target state is a finite GKP state,
@@
3.2 Gaussian operations
As we consider nonlinear dynamics, a relevant question is whether it is still possible to implement Gaussian operations within this framework. In our setting, this typically involves protocols that transiently generate non-Gaussian states during the evolution, but yield a final state that is Gaussian. A relevant example is the preparation of squeezed vacuum states, for which we demonstrate two distinct realizations. The first protocol employs only displacement modulation, while the second relies solely on depth modulation. Both cases are shown in Fig. 3 and achieve high final fidelities. The second example represents a generalized version of the frequency-jump protocol, in which a sudden change in potential depth excites breathing in the wave function, resulting in squeezing. In our case [Fig. 3(b)], the achieved squeezing surpasses what would be expected from a single-frequency-jump protocol, which is fundamentally limited by the ratio of initial and final frequencies.
Figure 3: Gaussian squeezing of the initial ground state via (a) displacement and (b) modulation. The potential is taken to be Gaussian with $\eta=0.15$. Protocols fidelities read: (a) $99.8\%$ (b) $99.6\%$. See the caption of Fig. 1 for subplot and curve legend details.
Figure 4: (a) Temporal evolution of a cat state in a perfectly harmonic potential. It rotates in the phase space and fully revives every trap period. (b) Evolution of a cat state in a tweezer characterized by $\eta=0.25$. (c) Evolution of a cat state in a tweezer characterized by $\eta=0.25$, however, with optimally controlled potential position and depth. (d) Fidelities between a cat state rotating in a perfectly harmonic potential and in Gaussian potentials with different nonharmonicities without (shades of green) and with (red) optimal control. See the caption of Fig. 1 for subplot and curve legend details.
3.3 State stabilization
Note that undriven evolution is nonlinear and, hence, the prepared state gets distorted during the evolution. In contrast, when evolved in the perfectly harmonic potential, the state undergoes rotation in the phase space with the trap period. In a realistic setup, such a state preservation in the rotating frame is destroyed by the nonharmonicity of the potential. The remedy lies in either decreasing the nonharmonicity or, alternatively, mimicking perfect phase space rotation through optimal control [60]. We present the latter method, which consists of two optimization steps. The first one involves using the average fidelity with respect to the perfectly rotating state during a multiple of the trap period $\tau_{\text{max}}=n2\pi$ as a reward function,
[…]
where
[…]
The second step guarantees high fidelity, should the protocol be applied consecutively many times, while the first one ensures that the state is not distorted at any time. In Fig. 4, we present an example of an optimally controlled cat state stabilization. The optimization is conducted for a single trap period, $\tau_{\text{max}}=2\pi$, with both position and depth control of a Gaussian potential. Subsequently, the optimized control is applied six times consecutively over six trap periods and compared with the bare evolution in a Gaussian potential. The results showcase that the presented optimal stabilization of the state surpasses that of a Gaussian potential with a relatively low nonharmonicity $\eta=0.05$, within this particular time scale.
Figure 5: (a,b) Two orthogonal four-legged-cat states spanning a two-level subspace evolved in a position- and depth-controlled single Gaussian well, realizing a $\sigma_{x}$ operation. (c) Cat state preparation using a double-well optimal control. The superscript L(R) signifies the state centered in the left(right) potential well. (d,e) State transfer between the wells—each of the vibrational states, $\psi_{0}(x)$ and $\psi_{1}(x)$, is transferred to the other well, without altering the relative phase. See the caption of Fig. 1 for subplot and curve legend details.